Hi,
Math problems I have no idea how to figure out and understand.I am not sure they are equation category. Please help
x-4/11=3/5
Divide & simplify-32/45+48/35
Simlify-6x*2-2y*2-5(3y*2+2x*2)
3(2x-1)=27-2
Clear fraction--3/2x1=2/5x4
Thank you so much, Betty

Answer by Boreal(15235)

x-(4/11)=(3/5)
multiply everything by 55
55x-20=33
55x=53
x=(53/55)
============
3(2x-1)=25
6x-3=25
6x=28
x=(14/3)
=========
(3/2x)=(2/5x^4)
cross multiply
15x^4=4x
x^3=(4/15)
x=(4/15)^(1/3)
=========
-32/45+48/35 ;;This has a common denominator of 1575. Multiply -32 by (35) {-1120} and 48 by 45 {2160}. The answer is 1040/1575, and that reduces to 208/315.
Simlify-6x^2-2y^2-5(3y^2+2x^2) I am editing out the asterisks to put in exponents. This may or may not be what was intended.
-6x^2-2y^2-15y^2-10x^2=
-16x^2-17y^2